OpenStudy (anonymous):

Lesley has a gift card worth $500 for a local entertainment store. Movies cost $20 each and newly released video games cost $50 each. In order to use the gift card, Lesley must purchase at least eleven items. Which is a combination of movies and newly released video games that Lesley could purchase using his gift card? 9 movies and 4 video games 3 movie and 10 video games 8 movies and 2 video games 13 movies and 6 video games

OpenStudy (shane_b):

Simply try each one: 9($20)+4($50)=$380 3($20)+10($50)=$560 8($20)+2($50)=$260 13($20)+6($50)=$560

OpenStudy (shane_b):

Two are possible answers...and two are not.